A car covers its outward journey at 40kmph. It is towed back, over exactly the same distance at 10mph. What is the average speed over the entire journey?
 
A)12mph B)14mph C)15mph D)20mph 

Grade:10

3 Answers

Arun
6 years ago
Dear student
 
we can apply the formula 2xy/(x+y) for average speed.
Hence
 
Avg speed = 2*40*10/(40+10) = 800/50 = 16 kmph
 
Regards
Arun (askIITians forum expert)
SUDHARSANA V V
6 years ago
Before working out Km is to be converted in to miles isn`t it. Hence 40km isto be converted and then average speed to be found out.
Arun
6 years ago
Sorry, thanks for pointing it out.
 
40 kmph = 40/1.6 mph = 24.85 mph
 
Now avg speed = 2*24.85*10/34.85 = 14.26 mph
 
Hence option B is correct